A stock has had returns of −19.3 percent, 29.3 percent, 27.6 percent, −10.4 percent, 35.1 percent, and 27.3 percent over the last six years.

What are the arithmetic and geometric returns for the stock? (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Arithmetic Returns for the Stock = Sum of returns / No. of returns

= (-19.3+29.3+27.6-10.4+35.1+27.3)/6

= 89.6/6

= 14.93%

Geometric Returns for the Stock = (R1*R2*.....*Rn)^(1/n)  

= (-.193*.293*.276*-.104*.351*.273)^(1/6)

= 0.00015553821^(1/6)

= 0.23190290167

= 23.19%